OpenStudy (anonymous):

A boat travels 3 hrs. with a current at a speed at 10kph and returns the same distance against the current an hour home. Determine the boat's speed in still water.

OpenStudy (anonymous):

when you say 'an hour home' what do you mean?

OpenStudy (anonymous):

by the way, it you return to the question about 2 numbers adding up to 114 i can explain how to do it

OpenStudy (anonymous):

yes, pls explain:))

OpenStudy (anonymous):

ok i'll go there now

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

right i've left an explanation in that post - though it took a while the system just went into crawl mode.

OpenStudy (anonymous):

let speed of boat = s speed of current = 10 kph d= distance travelled speed of boat along current s1 = s+10 time t1 = 3hours so distance traveled d1 = speed*time = (s+10)*3 = 3(s+10)=3s+30 -----(I) speed of boat against current s2 = s-10 time t2 = 4 hours so distance traveled d2 = speed*time = (s-10)*4 = 4(s-10)=4s-40 -----(II) but d1=d2 ( as same distance is traveled in both directions) so 3s+30 = 4s-40 so s=70 so the boat's speed in still water is 70 kph